The distance between two bodies is r , if the mass of one is doubled , so the distance between them has to be X*r to make attraction force between them decreases to its half.
1
Expert's answer
2019-03-22T10:37:43-0400

"\frac{m_1m_2}{r^2}\frac{1}{2} = \frac{2m_1m_2}{(Xr)^2} ; \\
X^2 = 4; \\
X = 2."